### v1

@@ Abstract
Machine learning (ML) is a rapidly developing area of medicine that uses significant resources to apply computer science and statistics to medical issues. ML’s proponents laud its capacity to handle vast, complicated, and erratic medical data. It’s common knowledge that attackers might cause misclassification by deliberately creating inputs for machine learning classifiers. Research on adversarial examples has been extensively conducted in the field of computer vision applications. Healthcare systems are thought to be highly difficult because of the security and life-or-death considerations they include, and performance accuracy is very important. Recent arguments have suggested that adversarial attacks could be made against medical image analysis (MedIA) technologies because of the accompanying technology infrastructure and powerful financial incentives. Since the diagnosis will be the basis for important decisions, it is essential to assess how strong medical DNN tasks are against adversarial attacks. Simple adversarial attacks have been taken into account in several earlier studies. However, DNNs are susceptible to more risky and realistic attacks. The present paper covers recent proposed adversarial attack strategies against DNNs for medical imaging as well as countermeasures. In this study, we review current techniques for adversarial imaging attacks, detections. It also encompasses various facets of these techniques and offers suggestions for the robustness of neural networks to be improved in the future.
@@
Adversarial attack, Medical image, Deep Neural Network, Model safety, Robustness
@@ 1 Introduction
1.1 Adversarial Attack
Adversarial attacks encompass a set of techniques aimed at manipulating machine learning models by introducing well-crafted, often imperceptible alterations to input data. These modifications aim to deceive the model, leading to misclassifications or inaccurate predictions [1]. Adversarial attacks have the potential to compromise machine learning systems’ dependability and security, which is concerning for crucial applications such as cybersecurity, autonomous vehicles, and medical diagnosis. Of all the adversarial attack types, ”White-box attacks” and ”Black-box attacks” are the most common.
Ongoing research in machine learning security centers on the analysis of adversarial attacks and the development of defense strategies. Researchers continually explore novel attack tactics to identify potential weaknesses in AI systems while concurrently striving to construct robust models less susceptible to adversarial perturbations. Artificial intelligence (AI) modern technology is widely successful in fields like computer vision, natural language processing, and automated driving; however, its application in vital safety sectors is hindered by its susceptibility to adversarial attacks. Consequently, enhancing the resilience of AI systems against such attacks has become paramount for the progress of AI [2].
The rapid advancement of AI technologies has found diverse applications in numerous domains, from machine translation, speech recognition, and object identification to more intricate tasks like drug composition analysis [3, 4, 5, 6, 7]. Noteworthy applications include brain circuit construction [8], particle accelerator data analysis [9], [10], and DNA mutation impact analysis [11]. Since Szegedy et al.’s seminal work [12] highlighting neural networks’ vulnerability to adversarial attacks, research on adversarial technologies for artificial intelligence has steadily expanded. New techniques for adversarial attacks and strategies for mitigating them are consistently emerging.
The term ”adversarial attacks at the training stage” refers to actions taken by adversaries to manipulate the training dataset, input characteristics, or data labels during the training phase of the target model. This manipulation of the training dataset involves actions such as adding or deleting training data, as demonstrated by Barreno et al.’s approach [13]. During the testing phase, adversarial assaults are categorized into white-box attacks and black-box attacks [2]. In white-box situations, attackers have access to the parameters, techniques, and structure of the target model, enabling them to craft adversarial samples based on this information.
@@ 1.2 The Adversary’s Objective: Evasion attack versus poisoning assault
”Poisoning attacks” are assault methods that allow an attacker to add or change a large number of fake samples to the training set of a DNN algorithm. The trained classifier may perform poorly as a result of these bogus data. They may have poor accuracy [14] conversely make imprecise predictions on a few samples for analysis [15]. The classifiers used in evasion attacks are fixed and often work well on safe testing samples. The adversaries are unable to reform the classifier’s settings or parameters, but they do generate some fictitious instances that the classifier is unable to distinguish.
@@ 1.3 Deep Neural Networks: Adversarial Attacks
Regarding medical image processing use cases such as diagnosing cancer and lesion identification, deep neural networks, also known as DNNs, are rapidly gaining popularity [16]. Nonetheless, a new study suggests that adversarial examples/attacks with tiny, barely noticeable changes can weaken medical deep learning systems. There are presently safety concerns associated with the usage of these devices in clinical settings [16]. Deep neural networks (DNN) are increasingly well-liked and effective at a variety of machine learning requisitions. They have been employed with surprising effectiveness in a variety of recognition issues in the fields of pictures, graphs, text, and voice [1]. They are able to identify things in images with accuracy that is almost human [17], [18]. Additionally, they are employed in speech recognition [19], natural language processing [20], and gaming [21].
@@
Regarding the data set provided by MNIST, Biggio et al.[22] produce adversarial instances first, focusing on traditional machine learning classifiers like SVMs and 3-layer neural networks with full connectivity which is shown in Figure 1. In order to trick the classifier, it optimizes the discriminant function. As an illustration, consider a linear SVM classifier on the MNIST dataset.
@@ 1.4 Examining examples of opposition in the real world
The research [23] put decals to traffic signs that pose a serious threat to autonomous vehicles’ sign recognition technology. These hostile objects are especially detrimental to deep learning systems since they can interfere directly with numerous real-world DNN usages, such facial recognition and autonomous cars. By assessing the adversarial images (FGSM, BIM) produced to see if they are ”robust” against changes in natural circumstances (e.g., shifting viewpoint, illumination, etc.), the authors of the work [24] investigate the viability of creating tangible adversarial objects. Robust in this context means that even after transformation, the produced pictures are still antagonistic. The experiment’s findings show that many of these adversarial examples—particularly those produced by FGSM—remain antagonistic to the classifier even after transformation. The findings imply that antagonistic real-world objects could trick the sensor in many situations.
@@ 1.5 Medical Image Under Artificial Intelligence Attack
Researchers have access to strong models of developing science and technology thanks to deep learning. Convolutional neural networks, also known as CNNs, are among the most significant categories of deep learning (DL) frameworks for the analysis and processing of images because of their exceptional ability to learn valuable information. The processing of the human organism utilizing different picture modalities for therapeutic, diagnostic, and health surveillance purposes is known as medical image analysis [25]. The advancement of computer vision through the use of deep neural networks addresses issues that were not well-solved by traditional image processing methods. Beinfeld et al. [26] asserted that \$385 spent on medical imaging results in a savings of almost \$3000. MRI, CT scans, ultrasound (US), and X-rays are the most frequently used image modalities. Figure 2 shows how adversarial assaults may be used to arbitrary alter diagnosis outcomes in three different medical picture datasets: Dermoscopy [27], X-ray of the chest [28], and Fundoscopy [29].
Figure 2: Instances of adversarial approaches designed by the PGD (Projected Gradient Descent) to deceive DNNs developed using datasets of medical images include dermoscopy [27] (third row), chest x-ray [28], and fundoscopy [29] (first row, DR=diabetic retinopathy). Normal images on the left, adversarial perturbations in the middle, and adversarial images on the right. The left-bottom tag indicates the projected class, and green or red indicates whether the predictions were accurate or not [16].
Hu et al. [1] assert that in order to develop more reliable models, it is crucial to investigate the reasons why adversarial cases emerge and to comprehend deep learning models better. Depending on the knowledge of the enemy, attacks can be categorized into three kinds. The saliency (or attention) map of a picture that was input shows the regions that significantly change the model’s output based on the gradients of the categorization loss with regard to the input [16]. We can see that some medical photos have highly concentrated regions that are noticeably larger. This could indicate that the rich biological textures included in medical photos occasionally entice the DNN model to focus more on aspects unrelated to the diagnosis. Tiny adjustments in these high-focus areas can significantly affect the model’s output.
@@
Byra et al. [30] proposed an attack strategy on ultrasonography (US) imaging for liver fatty tissue. Radio-frequency signals are used to rebuild US pictures, and the reconstruction technique was subjected to a zeroth-order optimization attack [31]. The InceptionResNetV2 model was used in the studies, and the assault resulted in a 48% loss in model accuracy. Ozbulak et al. [32] proposed an attack specifically designed for medical picture segmentation called the adaptive segmentation masking attack (ASMA). The suggested attack offers significant intersection-over-union (IoU) degradation and produces nearly undetectable samples for most portions. Because the U-Net framework is among the most well-known models for clinical picture segmentation, they employed it in the trials. The datasets used were for segmentation of glaucoma optic disk [33] and ISIC skin lesion segmentation [34]. Chen et al. published a method for fabricating hostile cases to thwart medical picture segmentation [35]. In order to simulate anatomical and intensity fluctuations, geometrical deformations are used to create the adversarial examples. By attempting to partition organs from abdominal CT scans using a U-Net model, they tested the effectiveness of these examples. With reference to the Dice score measure, they successfully reduced it significantly across all organs. The kidneys and pancreas, however, require a higher level of disturbance and are more arduous to assault than the liver as well as spleen.
@@
Tian et al. [36] examined the phenomena of bias field, which can arise from incorrectly capturing a medical image and jeopardize a DNN’s efficacy, as shown in Figure 3. Motivated by adversarial attacks, the authors created an adversarial-smooth bias discipline approach to trick a model. The ResNet50, MobileNet, among others and DenseNet121 models were employed to fine-tune the chest X-ray dataset that was used in their study. They looked at this attack’s transferability and white-box attacks. Comparing the suggested attack to other cutting-edge white-box attacks, it exhibited a greater attack accuracy on transferability.
@@ 2 Background Literature Review
Adversarial instances are intentionally created data inputs that are intended to degrade a machine learning model’s performance. When researchers examined strategies used by spammers to evade spam filters in 2004, they unofficially coined the term ”adversarial inputs” [37]. These misleading examples are usually produced by deliberately faking real data, such spam advertising messages, in order to trick the computer system that analyzes it. Alterations can be applied to text data, like spam, by introducing innocuous text or modifying frequently used phrases in malicious communications with synonyms. To fortify algorithms against adversarial attacks, researchers have explored various strategies, including training algorithms on adversarial samples and employing sophisticated data processing techniques to minimize the susceptibility to manipulation. The quest for fully robust models in machine learning aims to accelerate the development of algorithms capable of making decisions based on consistent explanations, with promising early efforts in this direction [38].
@@
In the realm of healthcare, medical claims codes play a crucial role in determining the amount spent on a patient visit post-payer approval. Payers often assess these claims using automated fraud detectors increasingly driven by machine learning. Historically, healthcare providers have shaped payer records of patient visits, including the associated codes, to influence the algorithmic outputs of payers [38]. Medical fraud, a market valued at \$250 billion, exemplifies the extreme end of this strategic tailoring of patient presentations. While some practitioners may overtly fabricate medical claims, patient data falsification often takes more covert forms. For instance, consistently submitting codes for billing services identical to, but more expensive than, those actually provided is known as intentional upcoding. Here, Table 1 shows past exclusive past research works where authors did some fantastic work for medical images to defend adversarial attacks. Different researcher proposed different frameworks to overcome this problem and to suggest a new thought for revolution.
2.1 Elements of a Cyber Adversary’s Assault
Figure 4: A showcase of executing adversarial attacks on various medical AI systems without resorting to overtly deceptive data manipulation [39].
In some instances, subtle adjustments to billing codes blur the boundary between fraudulent activities and conscientious best practices. One notable illustration may be found in the guidelines on the Endocrine Society website, which advise medical professionals not to bill for metabolic syndrome, or International Classification of Diseases (ICD) code 277.77, in situations of obesity. Coverage denial is likely to occur with this particular code/condition combo in Figure 4. The Society recommends an alternative approach, suggesting billing for the individual codes related to specific disorders comprising the metabolic syndrome, such as hypertension.
Table 1: Existing adversarial attack research on medical Image
Reference Year Framework Contribution
[40] 2021 Propose a framework to defend against adversarial training, speckle-noise attacks, and maintain accurate classification labels in recognizing diabetic retinopathy through analyzing retinal fundus images. a) accuracy of 99%; b) defensive model’s robustness ; c) a system that includes a fresh SN attack.
[41] 2022 Explain the use of gradient-free trained sign activation systems in medical imaging AI systems to identify and thwart hostile attacks. The model performs better and even twice as well as the finest.  a) Owing to significant distortion and excellent transferability accuracy, the model’s average success rate in adversarial case classification is 88.89%; b) This model defends against adversarial attacks.
[42] 2020 This method detects diverse adversarial attacks while preserving user anonymity and classification performance, making it a versatile addition for deep learning-based medical imaging systems to enhance resilience. a) The chest X-ray dataset exhibits robust performance across diverse configurations, contributing to the heightened safety of deep learning-based medical image classification systems; b) To improve medical image categorization systems based on deep learning security.
[43] 2022 In weakly-supervised clinical tasks, CNNs and ViTs exhibit vulnerability to both white- and black-box adversarial attacks with comparable baseline performance. a)Compared to CNNs, aViTs have a larger latent portrayal of clinically meaningful categories; b) Findings align with prior theoretical investigations and provide tangible evidence of ViTs’ capacity to grasp computational pathology; c) This suggests that computational pathology AI models will be deployed widely.
[…]
@@ 2.2 The components of an adversarial attack
Figure 5: Distribution of chosen articles on a yearly basis Pie chart; line graph [44].
Presenting a thorough overview of the various adversarial attack strategies and defense techniques is the aim of the study [44]. We first discuss the theoretical underpinnings, practices, and practical uses of adversarial attack strategies. Then, a brief discussion of the research on defense strategies obscuring the boundary of the large field follows. A few selected articles are published year-by-year from 2012 to 2021, as seen in Figure 5. This work aims to provide thorough taxonomies, evaluations of harmful assaults, and protections against the full DL pipeline. In this context, the numerous attack and defense tactics that have been developed over the previous two years have been categorized, with a focus on the clinical deep learning systems that are susceptible to adversarial attacks.
Several defensive strategies have been put forth in an effort to neutralize possible threats. A popular method in natural imaging called adversarial training adds adversarial images to the training dataset in order to make Convolutional neural network, or CNN, models more resilient. Nevertheless, this method is not the best for datasets pertaining to medical imaging, since the addition of different adversarial images could considerably reduce classification accuracy. A robust detection technique for malicious images is presented in the work of Li et al. [42], successfully defending against attacks on deep learning-driven medical image labeling systems. In a different study [43], scientists compared CNN performance with Vision Transformers (ViTs) to determine how resilient CNNs are to different types of attacks in computational pathology. The authors developed robust neural network models, evaluating their efficacy against both white- and black-box attacks. The structures of attacks for both models were scrutinized, with an exploration of the underlying factors influencing their performance. The study’s findings were validated through two clinically relevant classification tasks involving distinct patient groups. Preceding Ma et al.’s research [16], ambiguity surrounded medical image adversarial attacks (AAs), limiting adversarial machine learning analyses to natural images. Unlike natural images, medical images may contain domain-specific elements, impacting medical deep-learning systems susceptible to AAs. AAs possess the potential to manipulate diagnoses and outcomes, emphasizing the critical need for a robust healthcare infrastructure to mitigate potential risks from malicious attacks.

### v2

@@ Abstract
Machine learning (ML) is a rapidly developing area of medicine that uses significant resources to apply computer science and statistics to medical issues. ML’s proponents laud its capacity to handle vast, complicated, and erratic medical data. It is well known that adversaries can force misclassification by constructing inputs maliciously into machine learning classifiers. In the field of computer vision applications, such adversarial examples have been thoroughly researched. Healthcare systems are thought to be highly difficult because of the security and life-or-death considerations they include, and performance accuracy is very important. Recent arguments have suggested that adversarial attacks could be made against medical image analysis (MedIA) technologies because of the accompanying technology infrastructure and powerful financial incentives. Since the diagnosis will be the basis for important decisions, assessing how strong medical DNN tasks are against adversarial attacks is essential. Simple adversarial attacks have been taken into account in several earlier studies. However, DNNs are susceptible to more risky and realistic attacks. The present paper covers recent proposed adversarial attack strategies against DNNs for medical imaging as well as countermeasures. In this study, we review current techniques for adversarial imaging attacks, detections. It also encompasses various facets of these techniques and offers suggestions for the robustness of neural networks to be improved in the future.
@@
Adversarial attack; Medical image; Deep Neural Network; Model safety; Robustness
@@ 1 Introduction
In recent years, machine learning has been used as a transformative tool, particularly in medical image analysis. Particularly the processing ability of vast and complex medical data showed potential diagnostic accuracy and gave better outcomes. However, though there are advancements in machine learning security, there are several situations of adversarial attacks. These attacks, where inputs are maliciously altered to deceive the model, have been widely studied in computer vision but are less explored in the domain of medical imaging. This paper aims to explore recent adversarial attack strategies targeting DNNs in medical imaging, review detection techniques, and propose potential countermeasures to enhance the resilience and security of ML models in this critical field. We will go through the necessary sections to provide a critical review of these paper. The contributions of this research paper can be summarized below:
  1. 1.
    The research provides an in-depth overview of various adversarial attack strategies targeting deep neural networks (DNNs) in medical image analysis to identify vulnerabilities.
  2. 2.
    It examines existing DNN models used in medical image analysis and their resilience to adversarial threats.
  3. 3.
    The study evaluates the effectiveness of these DNN models against adversarial attacks.
  4. 4.
    Recommendations for enhancing security systems in medical image analysis are also provided.
1.1 Adversarial Attacks ML Applications
Adversarial attacks are a group of methods for altering machine learning models by adding expertly designed, frequently undetectable changes to input data. These alterations are intended to trick the model, causing it to misclassify data or forecast outcomes [1]. Adversarial assaults can endanger the dependability and security of machine learning systems, especially in crucial applications like autonomous vehicles, medical diagnosis, and cybersecurity. A number of adversarial attack types exist, but the two most popular subtypes are ”White-box attacks” and ”Black-box attacks.” An continuing area of research in the realm of machine learning security is the analysis of adversarial assaults and defense techniques. Researchers are always exploring novel attack tactics to find possible weaknesses in AI systems while also attempting to create robust models that are less vulnerable to adversarial perturbations. Computer vision, natural language processing, automated driving, and other industries have all benefited greatly from the application of artificial intelligence technology in recent years [2]. The uses of artificial intelligence (AI) technology in important security sectors are, however, constrained by the adversarial assaults that AI systems are susceptible to. Therefore, enhancing the resilience of AI systems against adversarial attacks has become more and more crucial to the advancement of AI [2]. Artificial intelligence technology applications have recently advanced quickly in many different industries. Artificial intelligence technologies have been used in fields such as image classification, object detection, voice control, machine translation, and more advanced ones like drug composition analysis [3], brain circuit reconstruction [4], particle accelerator data analysis [5], [6], and DNA mutation impact analysis [7] due to their high performance, availability, and intelligence. Since Szegedy et al.’s [8] hypothesis that neural networks are susceptible to adversarial attacks, research on artificial intelligence adversarial technologies has steadily become a hotspot, with new adversarial attack techniques and mitigation strategies being developed on a regular basis. The term ”adversarial attacks at the training stage” refers to actions taken by the adversaries to alter the training dataset, input characteristics, or data labels during the training phase of the target model. By changing or removing training data, Barreno et al.’s [9] alteration of the training dataset falls under the category of training dataset modification. Adversarial attacks can be categorized into white-box attacks and black-box attacks during the testing stage [2]. In white-box scenarios, the attackers have access to the target model’s parameters, methods, and structure. Adversaries can use this knowledge to create adversarial samples for attacks.
@@ 1.2 The Adversary’s Objective: Evasion attack versus poisoning assault
The term ”poisoning attacks” refers to the assaulting techniques that enable an attacker to insert/modify a number of fictitious samples into a DNN algorithm’s training database. The trained classifier may perform poorly as a result of these bogus data. They may have poor accuracy [10] or make incorrect predictions on some test samples [11]. The classifiers used in evasion attacks are fixed and often work well on safe testing samples. The adversaries are unable to alter the classifier’s settings or parameters, but they do create some fictitious samples that the classifier is unable to distinguish.
@@ 1.3 Deep Neural Networks: Adversarial Attacks
Deep neural networks (DNNs) have gained popularity for medical image processing applications like cancer diagnosis and lesion detection [12]. However, a recent study shows that adversarial examples/attacks with tiny, barely noticeable changes can weaken medical deep learning systems. The use of these devices in clinical settings is now accompanied by safety worries [12]. Deep neural networks (DNN) are increasingly well-liked and effective at a variety of machine learning applications. They have been used with surprising effectiveness in a variety of recognition issues in the fields of pictures, graphs, text, and voice [1]. They are able to identify things in images with accuracy that is almost human [13], [14]. Additionally, they are employed in speech recognition [15], natural language processing [16], and gaming [17].
@@
On the MNIST data set, Biggio et al.[18] produce adversarial instances first, focusing on traditional machine learning classifiers like SVMs and 3-layer fully-connected neural networks which is shown in Figure 1. In order to trick the classifier, it optimizes the discriminant function. As an illustration, consider a linear SVM classifier on the MNIST dataset.
@@ 1.4 Examining examples of opposition in the real world
The research [19] put decals to traffic signs that pose a serious threat to autonomous vehicles’ sign recognition technology. These hostile objects can directly interfere with many real-world DNN applications, such as face recognition, driverless vehicles, etc., making them more harmful to deep learning models. By determining if the created adversarial images (FGSM, BIM) are ”robust” under natural change (such as changing viewpoint, illumination, etc.), the authors of the work [20] investigate the viability of creating tangible adversarial objects. Robust in this context means that even after transformation, the produced pictures are still antagonistic. The experimental findings show that a significant fraction of these adversarial examples, particularly those produced by FGSM, continue to be adversarial to the classifier following transformation. The findings point to the possibility of actual hostile objects that can trick the sensor in various settings.
@@ 1.5 Medical Image Under Artificial Intelligence Attack
Researchers have access to strong models of developing science and technology thanks to deep learning. As they are highly good at learning useful features, convolutional neural networks (CNNs) are the most significant class of DL models for image processing and analysis. Medical image analysis is the processing of the human body using various picture modalities for diagnostic, therapeutic, and health monitoring purposes [21]. The development of deep neural networks in the field of computer vision addresses issues that were not well-solved by traditional image processing methods. Beinfeld et al. [22] asserted that \$385 spent on medical imaging results in a savings of almost \$3000. MRI, CT scans, ultrasound (US), and X-rays are the most frequently used image modalities. Figure 1 illustrates how diagnosis outcomes can be arbitrarily changed by adversarial attacks across three medical picture datasets: Fundoscopy [23], Chest X-Ray [24], and Dermoscopy [25].
Figure 2: Examples of adversarial approaches designed by the Projected Gradient Descent (PGD) to deceive DNNs trained on medical image datasets include dermoscopy [25] (third row), chest x-ray [24], and fundoscopy [23] (first row, DR=diabetic retinopathy). Normal images on the left, adversarial perturbations in the middle, and adversarial images on the right. The anticipated class is indicated by the left-bottom tag, and green or red denotes accurate or inaccurate predictions [12].
Hu et al. [1] assert that in order to develop more reliable models, it is crucial to investigate the reasons why adversarial cases emerge and to comprehend deep learning models better. Depending on the knowledge of the enemy, attacks can be categorized into three kinds. Based on the gradients of the classification loss with respect to the input, the saliency (or attention) map of an input image indicates the regions that significantly alter the model’s output [12]. We can see that some medical photos have highly concentrated regions that are noticeably larger. This could mean that the DNN model is occasionally drawn away by the rich biological textures present in medical images, causing it to focus more on details unrelated to the diagnosis. Small modifications in these areas of high focus can have a big impact on the model’s output.
@@
An attack method on ultrasound (US) imaging for fatty liver was put forth by Byra et al. [26]. Radio-frequency signals are used to rebuild US pictures, and the reconstruction technique was subjected to a zeroth-order optimization attack [27]. The InceptionResNetV2 model was used in the studies, and the assault resulted in a 48% loss in model accuracy. Adaptive segmentation mask attack (ASMA), a specific attack for medical picture segmentation, was suggested by Ozbulak et al. [28]. The suggested attack offers significant intersection-over-union (IoU) degradation and produces nearly undetectable samples for most portions. Because the U-Net model is one of the most well-known models for medical picture segmentation, they employed it in the trials. Datasets for ISIC skin lesion segmentation [29] and glaucoma optic disk segmentation [30] were employed. A technique for creating hostile instances to undermine medical picture segmentation was put out by Chen et al. [31]. In order to simulate anatomical and intensity fluctuations, geometrical deformations are used to create the adversarial examples. By attempting to partition organs from abdominal CT scans using a U-Net model, they tested the effectiveness of these examples. In terms of the Dice score metric, they successfully reduced it significantly across all organs. The kidneys and pancreas, however, require a higher level of disturbance and are more difficult to assault than the liver and spleen.
@@
As seen in Figure 3, Tian et al. [32] looked at the phenomenon of bias field, which can result from improperly acquiring a medical image and compromise a DNN’s effectiveness. The authors developed an adversarial-smooth bias field attack to deceive a model after being inspired by adversarial assaults. The chest X-ray dataset used for their research was tuned using the ResNet50, MobileNet, and DenseNet121 models. They looked at this attack’s transferability and white-box attacks. Comparing the suggested attack to other cutting-edge white-box attacks, it exhibited a greater attack accuracy on transferability.
@@ 2 Background Literature Review
Adversarial examples are data inputs that are specifically designed to make a machine learning model fall short. The term ”adversarial inputs” was first used informally in 2004, when academics looked into the strategies spammers used to get around spam filters [33]. Typically, threatening examples are created by intentionally altering genuine data, such as spam advertising messages, in order to deceive the algorithm that will process it. Such modifications can be made to text data, such as spam, by adding innocent text or by changing phrases that are frequently used in malicious communications to synonyms. Researchers have worked to create algorithms that are resistant to adversarial attacks, for example by exposing algorithms to hostile examples during training or utilizing cunning data processing to reduce the possibility of manipulation. Early efforts in this direction are encouraging, and we are hopeful that the pursuit of completely robust machine-learning models will spur the creation of algorithms that learn to make judgments for consistent justifications [34].
@@
Following payer approval, medical claims codes determine the amount paid for a patient visit. Payers often use automated fraud detectors, which are increasingly driven by machine learning, to assess these claims. Health care providers have historically shaped payers’ records of patient visits (and the related codes) to affect their decisions (the algorithmic outputs) [34]. Medical fraud, a \$250 billion market, is at the extremity of this tactical tailoring of a patient presentation. Although some practitioners might blatantly fabricate medical claims, patient data falsification frequently takes much more covert forms. Intentional upcoding, for instance, is the practice of routinely submitting billing codes for services that are similar to but more expensive than those that were actually rendered.
@@ 2.1 The components of an adversarial attack
The goal of the study [35] is to present a comprehensive overview of the various adversarial assault strategies and defense techniques. First, we talk about the theoretical foundations, procedures, and applications of adversarial attack tactics. The research on defense strategies covering the large field’s boundary is then briefly discussed. Figure LABEL:fig5 shows the year-by-year publishing of a few chosen articles from 2012 to 2021. In-depth taxonomies and analyses of harmful attacks and defenses against the full DL pipeline are the goals of this research. The numerous existing attack and defense strategies were categorized in this context, with a focus on those that have only been developed in the last two years, and the medical deep learning systems sensitive to adversarial attacks were examined.
Numerous defensive strategies have been suggested as a defense against the threat. Adversarial training, which enlarges the training dataset with adversarial images to improve the resilience of the trained Convolutional Neural Network (CNN) model, is a common technique in the natural imaging domain. This tactic is not ideal for medical imaging datasets, though, as the introduction of a large number of different hostile images into the training set can seriously reduce the classification accuracy. In article [36], authors suggest a reliable detection method for malicious images that can successfully fend off attacks on deep learning-based medical image categorization systems. In the study [37], scientists looked at how resilient CNNs in computational pathology were to various attacks and compared the findings to how resilient ViTs were. Authors also developed strong neural network models and assessed how well they defended against white- and black-box attacks. The attack structures for both models were examined, and the authors looked into why they performed as they did. The findings were verified by the authors using two clinically pertinent classification tasks on separate patient groups. Prior to the publication of Ma et al.’s study [12], medical image AAs were ambiguous and adversarial machine learning analysis was limited to natural image analysis. Medical images, in contrast to natural images, may include domain-specific elements. Medical deep-learning systems may be impacted by AAs. AAs have the discretion to alter diagnoses and results. Ma et al. emphasized the significance of the healthcare system’s extensive resource base. It invariably raises dangers where potential attackers might try to take advantage of the healthcare system.